### v2

# Improving Black-box Robustness with In-Context Rewriting

Abstract
Machine learning models often excel on in-distribution (ID) data but struggle with unseen out-of-distribution (OOD) inputs. Most techniques for improving OOD robustness are not applicable to settings where the model is effectively a black box, such as when the weights are frozen, retraining is costly, or the model is leveraged via an API. Test-time augmentation (TTA) is a simple post-hoc technique for improving robustness that sidesteps black-box constraints by aggregating predictions across multiple augmentations of the test input. TTA has seen limited use in NLP due to the challenge of generating effective natural language augmentations. In this work, we propose LLM-TTA, which uses LLM-generated augmentations as TTA’s augmentation function. LLM-TTA outperforms conventional augmentation functions across sentiment, toxicity, and news classification tasks for BERT and T5 models, with BERT’s OOD robustness improving by an average of 4.30 percentage points without regressing average ID performance. We explore selectively augmenting inputs based on prediction entropy to reduce the rate of expensive LLM augmentations, allowing us to maintain performance gains while reducing the average number of generated augmentations by 57.76%. LLM-TTA is agnostic to the task model architecture, does not require OOD labels, and is effective across low and high-resource settings. We share our data111https://huggingface.co/datasets/Kyle1668/LLM-TTA-Augmentation-Logs, models222https://huggingface.co/collections/Kyle1668/, and code333https://github.com/Kyle1668/LLM-TTA for reproducibility.
1 Introduction
Text classification models deployed in real-world settings must excel on in-distribution (ID) inputs sampled from their training distribution and be robust to unseen out-of-distribution (OOD) inputs. OOD robustness is important for deploying safe and trustworthy models in real-world settings (Hendrycks et al., 2021). This challenge is especially acute in high-stakes settings such as content moderation (Ashish et al., 2023; Zhou et al., 2021), spam detection (Dada et al., 2019), and healthcare (Rasmy et al., 2020). OOD robustness in NLP is challenging in practice due to the complex nature of natural language data, adversarial examples (Goyal et al., 2022), and shifting domains (Yuan et al., 2023; Koh et al., 2020; Yang et al., 2023).
Most existing methods for improving OOD robustness in NLP require access to model weights by modifying the training process (Howard & Ruder, 2018; Ma et al., 2019; Ruder et al., 2019; Tu et al., 2020; Yaghoobzadeh et al., 2021) or adapting the model to new domains at test time (Wang et al., 2023b). Modifying the task model can be challenging in practice when retraining is costly, the underlying model is abstracted away from the practitioner, or sufficient OOD labels are unavailable. These constraints render the task model effectively a black box. In this work, we study the NLP task of short-form text classification in a black-box setting by turning attention towards the inputs to the model.
Test-time augmentation (TTA) sidesteps the need for modifying the task model or new labels by aggregating multiple predictions over augmentations of the test input, thus arriving at more robust predictions. The choice of textual augmentation function is critical since augmentations must be diverse and semantic-preserving (Shanmugam et al., 2020), a challenge with conventional augmentation functions such as word insertion and synonym substitution (Xiong et al., 2023). LLM-driven advances in machine translation (Kocmi & Federmann, 2023; Wang et al., 2023a; Zhu et al., 2023), paraphrasing (Witteveen & Andrews, 2019; Wahle et al., 2022; Cegin et al., 2023), and style transfer (Patel et al., 2022; Suzgun et al., 2022; Roy et al., 2023) indicate that higher-quality augmentations are now feasible. In this work, we study using LLMs as the augmentation function for TTA (LLM-TTA).
We experiment with two LLM-TTA methods: zero-shot paraphrasing, where we prompt the LLM to generate paraphrases of the input text, and In-Context Rewriting (ICR), where the LLM rewrites the input to be more like a set of ID exemplars provided in the prompt. Both methods outperform TTA with conventional augmentation functions for BERT and T5 across averaged across nine datasets for sentiment, toxicity, and news topic classification. These results demonstrate that LLM-TTA is a simple black-box robustness technique effective across multiple tasks. Our primary findings are:
1.
LLM-TTA Improves OOD Robustness. ICR improves a BERT classifier’s absolute accuracy on OOD data by an average of 4.86% for sentiment, 6.85% for toxicity, and 1.18% for news topics, all with minimal regression to ID performance. For toxicity, BERT’s ID accuracy improves by 2.99%, suggesting that LLM-TTA can improve both ID and OOD performance in some settings. Ablating training set size (Section 5.4) shows that LLM-TTA is useful in data-scarce and rich settings.
2.
TTA with Conventional Augmentation Functions Often Hurts Performance. In contrast to LLM-TTA, TTA with conventional augmentation functions generally hurt both ID and OOD performance. Back-translation is the best-performing conventional augmentation functions, with BERT’s OOD robustness improving by an average of 2.85 percentage points while word insertion regresses performance by –0.22 points and substitution by –0.17 points averaged across tasks.
3.
Selectively Augmenting High-Entropy Test Inputs Improves Efficiency. We can reduce the rate of expensive LLM augmentations by only augmenting test inputs in which the task model is uncertain in its prediction. For BERT as the task model and ICR as the augmentation function, we reduce the percentage of test inputs requiring augmentation by an average of 57.76% while still improving robustness. The uncertainty threshold is only determined from ID statistics.
2 Related Work
Text Augmentation.
Data augmentation is a strategy that enables practitioners to significantly increase the diversity and quantity of data available without collecting additional annotations (Feng et al., 2021). Train-time augmentation has been found to improve performance in low-resource scenarios (Chen et al., 2021), mitigating harmful features such as gender bias (Zmigrod et al., 2019), and improving robustness (Morris et al., 2020; Ng et al., 2020).
Textual data augmentation can be performed at the character (Karpukhin et al., 2019), word (Zhang et al., 2015; Kobayashi, 2018; Wei & Zou, 2019), or whole-text (Vickrey & Koller, 2008; Hou et al., 2018; Yu et al., 2018; Anaby-Tavor et al., 2019; Kumar et al., 2020) level. One concrete example is word substitution (Wei & Zou, 2019), where each word in the text has some probability of being replaced with a related word. Increasing the likelihood of replacement can result in more diverse augmentations but comes with the risk of losing the original semantic meaning of the source example (Xie et al., 2019; Bayer et al., 2021). Conversely, augmentations that do not introduce sufficient diversity are unlikely to be effective for large pretrained models (Longpre et al., 2020).
Test-Time Adaptation.
This approach extends TTA by updating the weights of the source model at test-time (Liang et al., 2023). Adaptation is commonly implemented in an unsupervised manner by minimizing a proxy for supervised loss such as entropy (Wang et al., 2021; Zhang et al., 2021), distribution alignment (Hassan et al., 2023), or through perturbing model internals such as with dropout (Liang et al., 2023). Crucially, this technique assumes the ability to modify the source model. Singh & Ortega (2022) extended the NLP TTA evaluation from Lu et al. (2022) to include adaptation and found that entropy minimization-based adaptation can improve performance accuracy on OOD toxicity detection.
Test-Time Augmentation.
TTA refers to the practice of aggregating predictions across multiple augmentations of an original test input during inference. This technique has been widely used in computer vision to improve ID performance (Moshkov et al., 2019; Ashukha et al., 2021), OOD robustness (Enomoto et al., 2022; Kim et al., 2020; Molchanov et al., 2020), adversarial robustness (Song et al., 2017; Prakash et al., 2018; Cohen et al., 2019), and uncertainty estimation (Ayhan & Berens, 2018; Gawlikowski et al., 2021). Lu et al. (2022) provided the initial foray into studying TTA for NLP. They studied how common word-based augmentations improve the robustness of a distilBERT (Sanh et al., 2019) on the WILDS CivilComments (Koh et al., 2020) dataset. Matiana et al. (2021) combines cosine similarities of model representations across multiple paraphrases of the input to classify stories. Xiong et al. (2023) explored dynamically assigning different weights to augmentations during aggregation to reduce the influence of potentially noisy augmentations. These works found that incremental improvements in accuracy are possible but that the word-based augmentation functions are a bottleneck for performance. Our study differs from these works in that we focus our evaluation on LLM-based TTA and across a more diverse set of tasks, models, and augmentation functions, as well as measure ID performance alongside OOD robustness.
3 LLM-TTA: Generating Faithful Augmentations With LLMs
In this section, we describe LLM-TTA, a method for performing test-time augmentation for natural language classification using the rewriting capabilities of large language models.
3.1 Augmenting Out-of-Distribution Inputs
In this paper, we consider a supervised classification task from an input space to a discrete output space with classes. We assume access to a model trained on a dataset sampled from a distribution . At test-time, we are given a point sampled from an unknown distribution , which may or may not be the same as our original training distribution.
TTA is a simple post-training approach for improving the generalization accuracy of on arbitrary inputs . The main aim of TTA is to reduce the effects of a poor prediction on a single point by aggregating them across many similar augmented points. TTA involves three steps: augmentation, inferences, and aggregation. As described in Section 2, TTA differs from test-time adaptation (Wang et al., 2023b) in that the weights of the source model remain frozen. Unlike train-time augmentation (Shorten & Khoshgoftaar, 2019; Bayer et al., 2021), TTA does not require practitioners to modify their training regimes. Section 2 provides further description of these alternative techniques. We use terminology inspired by Shanmugam et al. (2020) and notation from Goodfellow et al. (2016).
1.
Augmentation. We define an augmentation as a stochastic transformation of an input from which we can sample another augmented point . Given a set of such augmentations , we define as a set of single samples from each of the augmentations in . For ease of notation, we assume one of the transformations in is the identity transformation such that the original point is always in .
2.
Inference. Each of the points in is passed into the model to generate a set of predictions
3.
Aggregation. A final prediction is derived from an aggregation function that combines the set of predictions . Common aggregation methods include vote-based aggregation, where the most commonly predicted class is chosen, and mean-based aggregation, where the probabilities for each class are first averaged across augmented samples before a prediction is made.
3.2 Rewriting OOD Inputs with Language Models
We study whether employing TTA with augmentations generated by an LLM (LLM-TTA) will improve a task-specific classifier’s robustness. LLMs have achieved strong performance on adjacent tasks that require faithfully rephrasing a target text. We hypothesize that the LLM-generated augmentations will outperform conventional augmentation functions. If so, these results suggest that LLM-TTA is a promising general non-parametric technique that NLP practitioners can leverage without depending on the task model architecture or access to weights.
In LLM-TTA, is an inference process where an augmentation of the natural language test input is generated by a language model (LLM) conditioned with a natural language prompt as well as the original input . We study two prompt templates for as detailed in Figure 3.
1.
Paraphrasing. contains instructions for the model to generate a semantic-preserving paraphrase of the test input. This approach is zero-shot in that contains no examples, thus not requiring access to ID data. We expect simple paraphrasing done by a capable LLM can generate diverse and semantic preserving augmentations, even for subtle labels such as sentiment.
2.
In-Context Rewriting (ICR). In this approach, we prompt the LLM to rewrite the given OOD example to be more like a set of ID exemplars. ICR does not require practitioners to articulate the semantics or writing style of the original distribution. Instead, the LLM is prompted to infer the differences between examples and via in-context learning. We hypothesize that rewriting OOD inputs to be more like ID data can further improve performance over simple paraphrasing since the task model was trained to excel on ID data.
3.3 Entropy-Based Selective Augmentation
LLM inference is inefficient when the model’s predictions are invariant across the test input and its augmentations. This invariance is a blocker for overruling an incorrect prediction for the test input. To reduce the rate of expensive LLM inferences, we explore whether the entropy of the task model’s class probability distribution is a predictive heuristic for whether the model will make an incorrect prediction. We can selectively augment test examples that are likely to benefit from TTA while deferring to the model’s original prediction otherwise. Lower entropy has been observed to be correlated with correct predictions in machine learning models (Grandvalet & Bengio, 2004). This observation has motivated entropy to be leveraged as an unsupervised loss metric in test-time adaptation techniques (Wang et al., 2021; Zhang et al., 2021; Singh & Ortega, 2022).
Entropy-based selective augmentation (ESA) can be formally defined as follows. Let be the test input, be the entropy of the output probability distribution of the model , and be an entropy threshold. For ease of notation, let be the aggregated prediction for using TTA with an augmentation function . We then define the model prediction as
TTA is only used when the original prediction entropy is above a predetermined threshold. The percentage of examples above this threshold (thus requiring augmentation) is referred to as the augmentation rate. We assume access to the labels for the ID evaluation set to determine an optimal threshold. Within this evaluation set, an optimal threshold is calculated that balances gains in accuracy while minimizing the augmentation rate, details of which are described in Appendix A.4.
4 Experimental Setup
We study how well LLM-TTA can improve the robustness of various task models across three short-form text classification tasks. We simulate a black-box setting by keeping the weights of the task model frozen at test time and study methods that are task model architecture independent. The following sections describe our datasets, task models, and TTA augmentation functions.
4.1 Datasets
We use three ID evaluation datasets and seven OOD datasets across sentiment analysis, toxicity detection, and news topic classification. Each task model is optimized for the ID evaluation set either through fine-tuning for BERT and T5, or prompting with Falcon. The training splits for the OOD datasets are not used. Dataset statistics are listed in Appendix A.6.
Sentiment Classification.
We consider three sentiment classification datasets from the BOSS benchmark (Yuan et al., 2023). Each is a three-way classification task where the labels are positive, neutral, and negative. The ID dataset consists of Amazon reviews (McAuley & Leskovec, 2013), while the three OOD datasets are DynaSent (Potts et al., 2020), SST-5 (Socher et al., 2013), and SemEval (Nakov et al., 2016). Yuan et al. (2023) selected these three OOD shifts since their centroids had low cosine similarity with the ID (Amazon) centroid.
Toxicity Detection.
We also leverage the toxicity task in the BOSS benchmark. Toxicity detection is framed as a binary classification task between non-toxic (negative) and toxic (positive). The ID dataset is Civil Comments (Borkan et al., 2019), a collection of users’ comments on articles on the Civil News platform. The OOD datasets are AdvCivil, an adversarial version of Civil Comments introduced in the benchmark, as well as existing datasets ToxiGen (Hartvigsen et al., 2022) and Implicit Hate (Elsherief et al., 2021). These shifts were similarly selected by Yuan et al. (2023) since they have low cosine similarity with the ID centroid.
News Topic Classification.
AG News (Zhang et al., 2015) is a four-class news topic classification problem where the model is tasked with determining whether a given news article pertains to “World," “Sports," “Business," or “Sci/Tech." This task has a single OOD dataset, a novel dataset that is composed of the AG News test set style-transferred to resemble social media posts. Since each entry in the ID evaluation set has a corresponding style transferred entry in the OOD set, we can isolate the effect that differences in writing style have on performance. Additional details on how this dataset was created can be found in Appendix A.8.
4.2 Task Models
We study black-box robustness for three task models. First, we finetune BERT (Devlin et al., 2019) and T5-Large (Raffel et al., 2020) for each ID dataset. Second, we consider Falcon-7b, an instruction-tuned LLM (Almazrouei et al., 2023). To make predictions with Falcon, we use 16-shot prompt with randomly selected ID exemplars with equal numbers per class. Finetuning details are in Appendix A.5, and Falcon prompting details are in Appendix A.10.
4.3 Baseline Augnmention Functions
We compare our LLM-TTA augmentation functions with three representative functions studied in the NLP literature.
Word-Level Augmentations.
Existing works studying TTA for NLP have focused on world-level augmentations (Lu et al., 2022; Xiong et al., 2023). We use the word insertion and substitution methods implemented in nlpaug recommended by Lu et al. (2022). We use this library’s default parameters, where each word in the text has a 30% chance of being augmented for a maximum of 10 words. Insertion adds a new word after a target word in the text based on the word BERT predicts will come next based on the preceding and subsequent words. Substitution follows a similar approach, with the target word being replaced. Using BERT’s predictions reduces the chance of adding nonsensical words that may change the semantics of the text.
Back-Translation.
We select EnglishGerman back-translation as a representative whole-text augmentation function. While not studied in the NLP TTA literature, translation is a common data augmentation technique (Edunov et al., 2018; Xie et al., 2019; Ng et al., 2019). Stochastic translations act as paraphrases of the original text. We select German as the target language since EnglishGerman is a common pairing in the literature (Edunov et al., 2018; Sennrich et al., 2015; Hoang et al., 2018). Including a whole-text augmentation function in our experiments allows us to study how well paraphrasing with and without an LLM affects performance. Model and decoding specifics are included in Appendix A.10.
4.4 TTA Settings
Aggregation.
Following Lu et al. (2022), each experiment uses the test input and four augmentations to arrive at a final prediction. We only use a single augmentation function per experiment rather than mixing augmentation functions to better study the effect that specific augmentation functions have on performance. We use two different aggregation methods for mapping the predictions over the augmentations to a final class prediction. For BERT, we average the class probability distribution across the test input and its four augmentations and select the class with the highest probability. T5 and Falcon use a vote-based aggregation method where a verbalizer function maps a set of task-specific valid tokens (such as “1", “pos," and “positive") to a class label. We then select the most commonly predicted class across the five inputs as the final prediction.
LLM-TTA.
We use Stable Beluga 2-7B (SB2) to generate augmentations. SB2 is a LLama 2 (Touvron et al., 2023) model fine-tuned with additional instruction tuning. Figure 3 details the prompts for the LLM-TTA paraphrasing and ICR methods. These prompts and the test input are passed to the model, resulting in four stochastic augmentations. Specifically, we use the HuggingFace generation API with four return sequences, four beam groups, four beams, and a diversity penalty of 0.5. The maximum number of new tokens is dynamically set to be four times the number of tokens in the test input. ICR uses 16 randomly selected unlabeled exemplars balanced across classes sourced from the ID training set.
ESA Threshold.
As introduced in Section 3.3, we study whether only augmenting test inputs in which the predicted class probability distribution is above a predetermined threshold still improves performance while reducing the augmentation rate. We find the optimal threshold for the ID test set that strikes a balance between performance gains and augmentation rate. We do not rely on access to OOD labels. We further describe this methodology in Appendix A.4.
5 Results
5.1 LLM-TTA Improves OOD Robustness
Generating augmentations with LLMs outperforms all other augmentation functions for BERT and T5 across each OOD task, as shown in Table 5.1. We report the average accuracy across the multiple OOD shifts for sentiment and toxicity. ICR is the best-performing augmentation function, with BERT’s OOD performance improving by an average of 4.30 percentage points and T5’s by 2.81 percentage points. Falcon benefits less with LLM-TTA regressing performance on sentiment and toxicity for an overall net regression of –3.01 percentage points.
ICR is the overall best-performing augmentation function. Augmenting test inputs to be more like ID exemplars outperforms 0-shot paraphrasing. These gains are achieved without requiring the entire ID dataset at test time, OOD labels, or explicit descriptions of the original distribution within the prompt. 0-shot paraphrasing also outperforms all conventional augmentation functions and ICR for the toxicity task. These results suggest that LLM-generated augmentations generally outperform conventional augmentations, and that performance can be further improved for some tasks by leveraging ICR over simple paraphrasing.
LLM-TTA can improve task model robustness without modifying the task model’s weights or training regime. TTA’s flexibility allows it to be slotted into existing systems. The degree to which TTA improves robustness is contingent on the augmentation function, with whole-text augmentation functions outperforming word-level augmentations. While we observe consistent improvements, more progress is needed to make models robust to OOD shifts as test time.
| Sentiment | Toxicity | News Tweets | |||||||
| Augmentation | BERT | T5 | Falcon | BERT | T5 | Falcon | BERT | T5 | Falcon |
| None | 52.05% | 57.99% | 47.16% | 53.90% | 58.81% | 66.68% | 88.57% | 89.01% | 25.96% |
[…]
| Substitute (Lu et al., 2022) | 51.78% | 55.07% | 44.27% | 53.37% | 57.65% | 58.78% | 88.86% | 89.13% | 24.88% |
| Translate (Sennrich et al., 2015) | 52.60% | 56.86% | 44.98% | 61.65% | 63.51% | 58.97% | 88.82% | 89.04% | 26.36% |
[…]
| LLM–TTA: ICR | 56.91% | 59.83% | 44.20% | 60.75% | 63.89% | 58.41% | 89.75% | 90.53% | 28.16% |
| Sentiment | Toxicity | News Tweets | |||||||
| Augmentation | BERT | T5 | Falcon | BERT | T5 | Falcon | BERT | T5 | Falcon |
| None | 90.38% | 90.12% | 90.49% | 88.46% | 90.57% | 17.01% | 94.43% | 94.87% | 30.25% |
[…]
| Substitute (Lu et al., 2022) | 89.96% | 85.78% | 86.58% | 89.90% | 91.33% | 19.99% | 94.86% | 91.91% | 26.34% |
| Translate (Sennrich et al., 2015) | 89.05% | 87.22% | 87.14% | 89.64% | 90.65% | 17.39% | 93.92% | 92.70% | 30.12% |
[…]
| LLM–TTA: ICR | 90.85% | 90.78% | 90.35% | 91.45% | 91.87% | 19.00% | 93.53% | 91.99% | 29.47% |
5.2 LLM-TTA Can Improve ID Performance
Preserving ID performance is essential in settings where the overall distribution has only partially shifted. Table 5.1 reports ID gains using TTA. LLM-TTA improves ID performance for BERT and T5 for sentiment and toxicity but not for news. ICR is the best-performing augmentation function for these models and distributions. All augmentation functions regress Falcon’s ID performance on sentiment and news. ID results follow a similar pattern to the OOD results reported in Table 5.1, where LLM-TTA performs the best for BERT and T5. These results suggest that LLM-TTA does not seriously regress ID performance and can even improve ID performance in some settings.
5.3 Selective Augmentation Improves Efficiency
| Sentiment | Toxicity | News Tweets | ||||
| Augmentation | Accuracy | Aug Rate | Accuracy | Aug Rate | Accuracy | Aug Rate |
| None | 52.05% | - | 53.90% | - | 88.57% | - |
[…]
| Paraphrase: ESA | 54.32% | 50.50% | 61.08% | 66.46% | 88.83% | 1.62% |
[…]
| ICR: ESA | 56.21% | 56.51% | 58.26% | 66.46% | 89.28% | 3.76% |
Tables 5.1 and 5.1 demonstrate that LLM-TTA can improve performance when every test input is augmented. However, augmenting each input can be costly when using LLMs. We introduced entropy-based selective augmentation (ESA) in Section 3.3 to address this issue. ESA only augments test inputs in which the model is uncertain in its prediction. We use the entropy of the OOD test input’s predicted class probability distribution as the confidence measure. The intuition behind this method is that TTA is unlikely to be effective when the task model is confident in its prediction for the unaugmented test inputs and is thus unlikely to change its prediction across faithful augmentations. We use BERT as our task model since this model explicitly outputs a class probability distribution and benefits the most from TTA.
Table 5.3 demonstrates that selective augmentation can preserve most performance gains while drastically reducing the number of expensive LLM calls. ICR tends to benefit more than paraphrasing. BERT’s robustness on sentiment is improved by 4.16 points when only augmenting 56.51% of inputs compared to 4.86 points when augmenting every input. Similarly, the majority of gains are still realized for toxicity while reducing the augmentation rate by roughly a third. News only preserved half of the performance gains yet reduced the augmentation rate by over 95% for ICR and paraphrasing.
That performance can still be improved while augmenting drastically fewer OOD inputs suggests that TTA does not change most model predictions. Selective augmentation is a promising direction for narrowing in on test inputs likely to benefit from augmentation while avoiding those that will not benefit. The entropy-based approach we studied can meet this criterion without requiring OOD labels. Future techniques can improve upon this approach by leveraging other signals beyond entropy.